11. Solve the following word problem (Hint: setup and solve a system of equations):
One number plus twice another number gives a sum of 42. The difference of these same two numbers equals –3.
Find the two numbers
a)The numbers are 12 and 15.
b)The numbers are 11 and 14.
c)The numbers are 13 and 16.
d)This problem has no solution.

Answer by stanbon(75887)

One number plus twice another number gives a sum of 42. The difference of these same two numbers equals –3.
Find the two numbers
----
Equations:
x + 2y = 42
x - y = -3
-----------------
Subtract and solve for "y":
3y = 45
y = 15
----
Solve for "x":
x - y = -3
x -15 = -3
x = 12
=================
Solution:
x = 12 ; y = 15
======================
Cheers,
Stan H.
---
a)The numbers are 12 and 15.
b)The numbers are 11 and 14.
c)The numbers are 13 and 16.
d)This problem has no solution.